OpenStudy (anonymous):

Curtis drank 1/2 of his glass of water. Isa drank 2 1/3 times more water than Curtis. How many glasses of water did Isa drink?

OpenStudy (anonymous):

basically, your problem is : 2 1/3 * 1/2 do you know how to multiply fractions ?

OpenStudy (anonymous):

idk

OpenStudy (skullpatrol):

write 2 1/3 as a single fraction

OpenStudy (anonymous):

2 1/3 is a mixed fraction...we will now turn it into an improper fraction. First, take the whole number and multiply it by the denominator. Then add that number to the numerator. Then put that number over the original denominator. 2 1/3 = (2 * 3 + 1) / 3 = 7/3 so our problem becomes : 7/3 * 1/2.....now just multiply straight across (7 * 1) / (3 * 2) = ??

OpenStudy (skullpatrol):

2 1/3 = 2.333...

OpenStudy (anonymous):

true ^^ you can also turn your fractions into decimals and just multiply them

OpenStudy (anonymous):

is the answer 1 1/6

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

7/6 = 1 1/6....yes, it is

OpenStudy (skullpatrol):

One and one-sixth glasses of water